OpenStudy (anonymous):

Factor completely 2xsquared +7x+4

OpenStudy (anonymous):

Is this a test question?

OpenStudy (anonymous):

no its for some hw

OpenStudy (anonymous):

In any Quadratic with three terms in the form of ax^2 + bx + c, take a * c. Once you have foud this product, see what numbers would be factors of a*c, and yet add up to b. These could be negative too. In your example, we will take 2 * 4 which is what?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

oh i messed up its -4

OpenStudy (anonymous):

Good, now what are some factors of 8 that when added together make 7?

OpenStudy (anonymous):

in the equation

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

ah! ok, so -8, so what are some factors of -8 that could add up to +7?

OpenStudy (anonymous):

OK, name some factors of -8.

OpenStudy (anonymous):

4,2 and 8

OpenStudy (anonymous):

OK, so -2 and 4 multiplied = -8, when we add them together we get +2, so this is not the one we are looking for. Any other factors?

OpenStudy (anonymous):

1 and other than that i have no idea

OpenStudy (anonymous):

ok, 1 * -8 = -8 and 1 + -8 = -7...close! Could we somehow change the sign to make these work?

OpenStudy (anonymous):

im guessing we can but im not sure how

OpenStudy (anonymous):

well if we tried -8 and +1, could we flip the signs?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So if we tried -1 and +8, then -1 * 8 = -8 and -1 + 8 = +7... it works!

OpenStudy (anonymous):

okay yay ahh then what do i do

OpenStudy (anonymous):

You need to be able to try all the factors and all combinations of positive and negatives.

OpenStudy (anonymous):

After that, you rewrite your original equation while using these two new numbers

OpenStudy (anonymous):

2x^2 -1x + 8x - 4

OpenStudy (anonymous):

Notice that we have not changed anything since -1x + 8x = 7x

OpenStudy (anonymous):

okay yeah

OpenStudy (anonymous):

from here, we will factor by grouping, so we will put parantheses around the first two terms, and then parantheses around the second two terms... (2x^2 - 1x) + (8x - 4)

OpenStudy (anonymous):

Now look at the first group, do these terms have anything that can be pulled out?

OpenStudy (anonymous):

factor but not the x

OpenStudy (anonymous):

Well since they both share an x, we can pull an x out!

OpenStudy (anonymous):

oh i was looking at the second one okay

OpenStudy (anonymous):

So 2x^2 - 1x = x (2x - 1)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

OK, now for the second group. Do they share anything?

OpenStudy (anonymous):

a common factor

OpenStudy (anonymous):

yes, what common factor?

OpenStudy (anonymous):

2 or 4

OpenStudy (anonymous):

and

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

4 is the largest common factor!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so 8x - 4 = 4(2x - 1)

OpenStudy (anonymous):

So now our equation 2x^2 + 7x - 4 = x(2x - 1) + 4(2x - 1)

OpenStudy (anonymous):

so now both terms have what in common?

OpenStudy (anonymous):

the 2x -1

OpenStudy (anonymous):

Excellent! So we factor that out and we get (2x - 1) (x + 4)

OpenStudy (anonymous):

Done!

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem!